A company is designing a product layout for a new product. It plans to use this production line eight hours a day in order to meet a schedule of 400 units per day. The tasks necessary to produce this product are detailed in the table below.

Task Predecessor A Time (seconds) 50 20 26 B A с А D 40 B.C C 30 m

a. Draw the network described in the table. 
b. What is the required cycle time (in seconds) in order to meet the schedule? 
c. What is the theoretical minimum number of workstations needed to meet the schedule? 
d. Balance this line using longest processing time. 
e. What is the efficiency of the balance obtained in part d?

please show work
Answer #1

CYCLE TIME = OPERATING TIME / DESIRED OUTPUT

AVAILABLE OPERATING TIME = 8 * 60 * 60 = 28800

CYCLE TIME = 28800 / 400 = 72

THEORETICAL NUMBER OF WORKSTATIONS = SIGMA(TASK TIME) / N NUMBER OF TASKS, N = 5

N = (50 + 20 + 26 + 40 + 30) / 72 = 3


WORKSTATION

ASSIGN TASK

TASK TIME

TIME REMAINING

IDLE TIME

1

A

50

72 - 50 = 22

B

20

22 - 20 = 2

2

2

C

26

72 - 26 = 46

D

40

46 - 40 = 6

6

3

E

30

72 - 30 = 42

42

CUMULATIVE IDLE TIME = 2 + 6 + 42 = 50

EFFICIENCY = (1 - (SIGMA(IDLE TIME) / (NUMBER OF WORKSTATIONS * CYCLE TIME))) * 100 = 1 - (50 / (3 * 72))) * 100 = 76.8519 OR 76.9%

B D A с E